Early Payment discount decisions:

Prairie Manufacturing has four possible​ suppliers, all of which offer different credit terms. Except for the differences in credit​terms, their products and services are virtually identical. The credit terms offered by these suppliers are shown in the following​table:

SUPPLIER        CREDIT TERM

J                      3/5        NET 30 EOM

K                      4/30      NET 100 EOM

L                      2/15      NET 60 EOM

M                     2/10      NET 120 EOM

.

​ (Assume a​ 365-day year.)

A. Calculate the approximate cost of giving up the cash discount from each supplier.

B. If the firm needs​ short-term funds, which are currently available from its commercial bank at

99​%, and if each of the suppliers is viewed separately​, which, if​ any, of the​ suppliers' cash discounts should the firm give​up?  

C. Now assume that the firm could stretch by 30 days its accounts payable​ (net period​ only) from supplier M. What​ impact, if​ any, would that have on your answer in part b relative to this​supplier?
Answer #1

A. Cost of giving up cash discount
J = 3%/97% x (365/25) = 45.15%
K = 4%/96% x (365/70) = 21.72%
L = 2%/98% x (365/45) = 16.55%
M = 2%/98% x (365/110) = 6.77%

B. Firm should give up cash discount of Supplier M, since it is lower than bank borrowing rate

C. Cost of Giving up
M = 2%/98% x (365/140) = 5.32%

There would be no impact on answer given in part b since stretching credit period will be lower the cost of giving up